# Enhancing ECG Classification Robustness with Lightweight Unsupervised Anomaly Detection Filters

### v1

@@ Abstract
Continuous electrocardiogram (ECG) monitoring via wearables offers significant potential for early cardiovascular disease (CVD) detection. However, deploying deep learning models for automated analysis in resource-constrained environments faces reliability challenges due to inevitable Out-of-Distribution (OOD) data. OOD inputs, such as unseen pathologies or noise-corrupted signals, often cause erroneous, high-confidence predictions by standard classifiers, compromising patient safety. Existing OOD detection methods either neglect computational constraints or address noise and unseen classes separately. This paper explores Unsupervised Anomaly Detection (UAD) as an independent, upstream filtering mechanism to improve robustness. We benchmark six UAD approaches, including Deep SVDD, reconstruction-based models, Masked Anomaly Detection, normalizing flows, and diffusion models, optimized via Neural Architecture Search (NAS) under strict resource constraints (at most 512k parameters). Evaluation on PTB-XL and BUT QDB datasets assessed detection of OOD CVD classes and signals unsuitable for analysis due to noise. Results show Deep SVDD consistently achieves the best trade-off between detection and efficiency. In a realistic deployment simulation, integrating the optimized Deep SVDD filter with a diagnostic classifier improved accuracy by up to 21 percentage points over a classifier-only baseline. This study demonstrates that optimized UAD filters can safeguard automated ECG analysis, enabling safer, more reliable continuous cardiovascular monitoring on wearables.
@@ I Introduction
Cardiovascular diseases (CVDs) are the leading cause of mortality worldwide, accounting for 19.8 million deaths in 2022 [1]. Early detection and timely intervention, encompassing pharmacological therapy and lifestyle modifications, are critical for improving patient outcomes and mitigating the associated healthcare burden. However, traditional monitoring within clinical settings often fails to capture the sporadic or asymptomatic cardiac events that may serve as early indicators of underlying pathologies. The advent of wearable technology offers a promising avenue for continuous, ambulatory electrocardiogram (ECG) monitoring, facilitating the detection of transient abnormalities without impeding patients’ daily activities. The large-scale deployment of continuous ECG monitoring generates vast quantities of data that preclude manual analysis, necessitating automated interpretation. Over the past decade, machine learning (ML), and particularly deep learning (DL), has demonstrated remarkable success in analyzing complex ECG signals [2].
Despite these advancements, the transition from controlled datasets to real-world clinical deployment, especially on resource-constrained wearable devices, poses significant challenges. Beyond the fundamental requirements of high accuracy and computational efficiency, a critical challenge arises from the inevitability of encountering Out-of-Distribution (OOD) data. In the context of ambulatory ECG analysis, OOD data typically manifests in two critical forms: ECG signals exhibiting pathologies unseen during the model’s training (unseen CVD classes), and signals that are so severely corrupted by noise artifacts (e.g., electrode motion or muscle activity) that they are rendered unsuitable for analysis. Standard classifiers are prone to making erroneous, often high-confidence, predictions on such inputs, which can undermine the reliability of the system [3, 4].
Addressing this challenge through purely supervised learning is impractical, as it requires an exhaustive, labeled dataset encompassing all possible CVDs and noise profiles encountered in daily life. Consequently, research has focused on methods for OOD detection and Uncertainty Quantification (UQ). Existing approaches often rely on analyzing the outputs or latent features of supervised classifiers (e.g., using MC Dropout, Deep Ensembles, post-hoc energy scores, or specialized architectures) [5, 6, 7, 8]. While these methods are valuable, they frequently address the detection of unseen classes and the identification of noise in isolation. Furthermore, many sophisticated approaches overlook the stringent computational constraints required for real-time deployment on wearable devices.
To address these gaps, this paper investigates the application of Unsupervised Anomaly Detection (UAD) as an independent filtering mechanism deployed upstream of a diagnostic classifier, as illustrated in Figure 1. UAD methods learn a compact representation of the in-distribution data without requiring labeled examples of anomalies, enabling the detection of significant deviations caused by either novel pathologies or noise. We conduct a systematic evaluation of diverse UAD methodologies, including Deep Support Vector Data Description (Deep SVDD), reconstruction-based methods (autoencoders (AE) and variational autoencoders (VAE)), Masked Anomaly Detection (MAD), Denoising Diffusion Probabilistic Models (DDPM), and normalizing flows (NF). Our analysis focuses specifically on the trade-off between detection performance and computational efficiency essential for deployment in resource-constrained environments.
@@
The main contributions of this work are summarized as follows:
  • We benchmark six distinct UAD methodologies, optimized via Neural Architecture Search (NAS) under strict resource constraints ($\leq$512k parameters). We evaluate their efficacy for the detection of OOD CVD classes and signals unsuitable for analysis due to noise using the PTB-XL and BUT QDB datasets.
  • We empirically identify Deep SVDD as the superior methodology for this application, demonstrating that it offers the optimal balance between computational efficiency and detection performance
  • We validate the efficacy of an integrated classifier-filter system in a realistic deployment simulation incorporating unseen CVD classes and calibrated real-world ambulatory noise. We show that the upstream UAD filter significantly enhances diagnostic robustness, yielding an accuracy improvement of up to 21.0 percentage points over a classifier-only baseline.
The remainder of this paper is organized as follows: Section II reviews related work. Section III details the datasets, the investigated UAD methodologies, and the experimental setup. Section IV presents the empirical results. Section V discusses the findings. Finally, Section VI concludes the paper.
@@ II Related Work
The deployment of ML models for ECG analysis in real-world clinical settings introduces significant challenges beyond controlled environments. Key among these are generalization across different data distributions (domain shift), robustness to varying signal quality, and the critical safety requirement of identifying inputs that fall outside the model’s training, known as OOD detection or open-set recognition.
II-A Domain Generalization and Distribution Shifts
A significant challenge in ML-based ECG analysis is performance degradation due to distribution shifts (e.g., variations in demographics or hardware). This is typically addressed as a domain generalization problem, aiming to maintain performance on known classes despite these shifts. Ballas and Diou [9] leveraged intermediate features from a ResNet-18 to improve generalization for 12-lead ECG classification across hospital databases. Similarly, Soltanieh et al. [10] demonstrated that Self-Supervised Learning (SSL) techniques (SimCLR, BYOL, SwAV) can learn representations that generalize well for arrhythmia detection across different OOD datasets. While these works address robustness to domain shifts, they do not specifically tackle the detection of entirely unseen disease classes or the identification of data unsuitable for analysis due to noise.
II-B Signal Quality Detection and Uncertainty Quantification
The reliability of automated ECG analysis is heavily dependent on input signal quality, as noise artifacts (e.g., muscle activity, electrode motion) prevalent in ambulatory monitoring can severely impair diagnostic accuracy. The literature addresses this through Signal Quality Assessment (SQA) and Uncertainty Quantification (UQ). Traditional SQA approaches focus on the explicit detection and classification of noise contamination. Satija et al. [11] proposed a framework using modified ensemble empirical mode decomposition (CEEMD) and temporal features to categorize ECG signals by noise type, facilitating targeted denoising or false alarm reduction. Alternatively, UQ methods address signal quality indirectly by identifying when a diagnostic model’s confidence is compromised due to impaired input. Jahmunah et al. [12] developed a Dirichlet DenseNet model to quantify uncertainty in myocardial infarction diagnosis. Utilizing predictive entropy, they demonstrated the model’s ability to convey low confidence when analyzing noisy signals, aiming to mitigate the risk of erroneous predictions based on corrupted data.
In contrast, our work employs UAD as an independent, upstream mechanism to explicitly filter out signals deemed entirely unsuitable for analysis due to noise, irrespective of the specific noise type, thereby preventing unreliable data from reaching the classifier.
II-C Unsupervised Anomaly Detection in ECG
@@
UAD methods identify deviations from a learned distribution of normal data without labeled anomalies. In ECG analysis, UAD has been explored primarily to differentiate abnormal patterns from normal sinus rhythm or to enhance supervised learning. Atamny et al. [13] conducted a comparative study of various unsupervised models, including AEs, VAEs, diffusion models, NFs, and Gaussian Mixture Models (GMMs). Trained solely on normal data, these models utilized reconstruction errors or likelihood estimations to differentiate abnormal ECG signals from healthy ones, with VAEs demonstrating the highest performance. In a different approach aimed at addressing the challenges of highly imbalanced (long-tail) datasets, Jiang et al. [14] employed self-supervised anomaly detection as a pretraining mechanism. They utilized masking and restoration techniques with multi-scale cross-attention to enhance feature representation, thereby improving the subsequent classification of rare cardiac anomalies.
While these works utilize UAD methodologies, their objectives differ significantly from ours. They focus respectively on binary diagnostic separation (healthy vs. abnormal) or improving the recognition of rare, but known, classes. In contrast, our work concentrates on deploying UAD as an independent, upstream filter designed to reject entirely unseen disease classes and signals rendered unsuitable for further analysis by noise.
II-D Open-Set Recognition in ECG Applications
Open-set recognition, the detection of unknown inputs during deployment, is also prevalent in biometric authentication using cardiac signals. Studies in this domain focus on rejecting unregistered subjects. Dong et al. [15] proposed a robust authentication system using multi-modal pretraining and specialized loss functions to enforce clear decision boundaries. Wu et al. [16] utilized compressed CNNs and user-specific feature vectors for efficient open-set identification. Although these studies address the open-set problem, the application domain involves fundamentally different objectives and data characteristics compared to clinical diagnosis.
II-E Detection of Unseen Cardiovascular Diseases
@@ III-A Datasets and Preprocessing
We used three datasets in this study. PTB-XL is a large, publicly accessible clinical ECG database comprising 21,799 12-lead ECG recordings, each 10 seconds in duration, from 18,869 patients in its most recent version 1.0.3 [17, 18, 19]. The recordings are annotated according to the SCP-ECG standard, covering diagnostic, form, and rhythm statements provided by cardiologists. Diagnostic labels are organized hierarchically into five superclasses: Normal (NORM), Myocardial Infarction (MI), Conduction Disturbance (CD), ST/T Change (STTC), and Hypertrophy (HYP). There are potentially multiple superclasses per record, as they are not mutually exclusive. The dataset facilitates standardized evaluation through pre-defined, patient-aware, 10-fold cross-validation splits, stratified by diagnosis, sex, and age. ECG signals are provided at both 500 Hz and 100 Hz sampling frequencies. We chose the 500 Hz samples.
The Brno University of Technology ECG Quality Database (BUT QDB) is designed for evaluating signal quality algorithms [20, 19]. It contains 18 long-term, single-lead ECG recordings (sampled at 1,000 Hz) collected during normal daily activities. Expert annotations classify signal quality into three categories: Class 1 (clear visibility of QRS complex, P waves and T waves), Class 2 (only QRS complexes reliably detectable), and Class 3 (signal unsuitable for analysis, QRS complexes undetectable).
The MIT-BIH Noise Stress Test database provides realistic noise recordings characteristic of ambulatory ECGs [21, 19]. It includes three half-hour recordings representing Baseline Wander (bw), Muscle Artifact (ma), and Electrode Motion Artifact (em). The ’em’ record, utilized in this study for noise injection, contains significant electrode motion artifacts along with baseline wander and muscle noise.
All ECG recordings were segmented into 10-second windows. These windows were resampled to a uniform length of 512 timesteps, a configuration determined during pre-screening tests to optimize the balance between computational efficiency and model performance. For the PTB-XL dataset, z-score normalization was applied, as it yielded superior performance. For the BUT QDB dataset, instance normalization proved more effective.
@@ III-B Unsupervised Anomaly Detection Methods
We investigated six UAD methods for detecting anomalous ECG signals. To optimize hyperparameters and evaluate the performance-efficiency trade-off, a NAS using random search (100 trials) was conducted for each method. Models were constrained to a maximum of 512k parameters to ensure suitability for resource-constrained devices.
Deep SVDD aims to map normal data into a compact hypersphere of minimum volume [22]. The training objective minimizes the distance of representations from a predefined center, enforcing constraints (no bias terms, unbounded activations, fixed center) to prevent hypersphere collapse. The anomaly score is the squared Euclidean distance to the center in the output space. A 1D ResNet architecture was employed. The NAS optimized the number of layers, filters, kernel sizes, strides, and the latent dimension size.
@@
MAD utilizes a self-supervised approach similar to masked autoencoders [23]. During training, 5% of time steps (determined via pre-screening) are masked with random values, and the model reconstructs the original values using bidirectional context. The anomaly score is the sum of reconstruction errors calculated by sequentially masking and reconstructing each time step during inference. A standard transformer architecture with an initial patching layer was used. The NAS varied the number of transformer blocks, hidden dimensions, attention heads, linear layer sizes, patch size, and patch overlap.
@@
Reconstruction-based methods (AE and VAE) assume that models trained solely on normal data will fail to accurately reconstruct anomalous inputs [24]. We utilized the L2 reconstruction error as the anomaly score. The VAE differs by employing a probabilistic encoder regularized by a KL divergence loss toward a standard normal distribution [25]. Stacked 1D convolution layers were used for both encoder and decoder architectures. The NAS optimized the number of layers, filters, kernel sizes, strides, and latent dimension size. For the VAE, the KL loss weight was also varied.
@@
DDPMs involve a fixed forward diffusion process that gradually adds Gaussian noise over T timesteps, and a learned reverse denoising process [26]. For anomaly detection, we adopted the approach of Wyatt et al. [27]: an anomalous input is partially diffused and then reconstructed. The model, trained only on the normal data manifold, attempts to ”repair” abnormal regions during denoising. A 1D U-Net architecture with convolution and attention layers was implemented using the denoising-diffusion-pytorch package [28]. The NAS varied the diffusion steps, U-Net architecture parameters (layers, filters, attention heads, hidden dimensions), and the U-Net objective (predicting noise, the original input or a variable from angular parametrization as in Salimans and Ho [29]).
@@
NFs are exact-likelihood generative models that learn complex data distributions by transforming a simple base distribution through invertible mappings [30, 31]. We implemented a multiscale GLOW architecture [32] with affine coupling layers and invertible 1x1 convolutions, utilizing the normflows package [33]. The exact probability density of a sample serves as the anomaly score. Stacked 1D convolutions were used for the subnetwork within the coupling layers. The NAS optimized the subnetwork architecture (layers, filters, kernel sizes) and the flow architecture (number of coupling blocks, split/squeeze dimensions and ratios).
Figure 2: Visualization of the realistic noise injection process applied to the PTB-XL dataset. A clean ECG signal (bottom left) is corrupted by adding calibrated noise (top center), sourced from the MIT-BIH noise stress test database. The resulting noisy ECG signal (bottom right) simulates real-world ambulatory conditions where the signal quality is unsuitable for analysis.

### v2

@@ Abstract
Continuous electrocardiogram (ECG) monitoring via wearable devices is vital for early cardiovascular disease detection. However, deploying deep learning models on resource-constrained microcontrollers faces reliability challenges, particularly from Out-of-Distribution (OOD) pathologies and noise. Standard classifiers often yield high-confidence errors on such data. Existing OOD detection methods either neglect computational constraints or address noise and unseen classes separately. This paper investigates Unsupervised Anomaly Detection (UAD) as a lightweight, upstream filtering mechanism. We perform a Neural Architecture Search (NAS) on six UAD approaches, including Deep Support Vector Data Description (Deep SVDD), input reconstruction with (Variational-)Autoencoders (AE/VAE), Masked Anomaly Detection (MAD), Normalizing Flows (NFs) and Denoising Diffusion Probabilistic Models (DDPM) under strict hardware constraints ($\leq$512k parameters), suitable for microcontrollers. Evaluating on the PTB-XL and BUT QDB datasets, we demonstrate that a NAS-optimized Deep SVDD offers the superior Pareto efficiency between detection performance and model size. In a simulated deployment, this lightweight filter improves the accuracy of a diagnostic classifier by up to 21.0 percentage points, demonstrating that optimized UAD filters can safeguard ECG analysis on wearables.
@@ I Introduction
The transition from controlled datasets to real-world clinical deployment, especially on resource-constrained wearable devices, poses significant challenges. Beyond the fundamental requirements of high accuracy and computational efficiency, a critical challenge arises from the inevitability of encountering Out-of-Distribution (OOD) data. In the context of ambulatory ECG analysis, OOD data typically manifests in two critical forms: ECG signals exhibiting pathologies unseen during the model’s training (unseen CVD classes), and signals that are so severely corrupted by noise artifacts (e.g., electrode motion or muscle activity) that they are rendered unsuitable for analysis. Standard classifiers are prone to making erroneous, often high-confidence, predictions on such inputs, which can undermine the reliability of the system [1].
@@
Addressing this challenge through purely supervised learning is impractical, as it requires an exhaustive, labeled dataset encompassing all possible CVDs and noise profiles encountered in daily life. Consequently, research has focused on methods for OOD detection. Existing approaches like uncertainty quantification (UQ) (e.g., using MC Dropout or Deep Ensembles) [2] can detect OOD data but suffer from inefficiency due to multiple ensemble members of forward passes. Furthermore, they frequently address the detection of unseen classes and the identification of noise in isolation.
To address these gaps, this paper investigates the application of Unsupervised Anomaly Detection (UAD) as an independent filtering mechanism deployed upstream of a diagnostic classifier, as illustrated in Figure 1. Our approach represents a single-inference architecture constrained to under 512k parameters, ensuring compatibility with the limited flash memory and SRAM of low-power microcontrollers. Since the filter can run independently from the classifier, the classifier will not be used when inputs are filtered out, further saving energy. UAD methods learn a compact representation of the in-distribution data without requiring labeled examples of anomalies, enabling the detection of significant deviations caused by either novel pathologies or noise. We conduct a systematic evaluation of diverse UAD methodologies, including Deep Support Vector Data Description (Deep SVDD), reconstruction-based methods (autoencoders (AE) and variational autoencoders (VAE)), Masked Anomaly Detection (MAD), Denoising Diffusion Probabilistic Models (DDPM), and normalizing flows (NF). Our analysis focuses specifically on the trade-off between detection performance and computational efficiency essential for deployment in resource-constrained environments. Our contributions are as follows: (i) We benchmark six distinct UAD methodologies, optimized via Neural Architecture Search (NAS) under strict resource constraints ($\leq$512k parameters) and evaluate their efficacy for the detection of OOD CVD classes and signals unsuitable for analysis due to noise using the PTB-XL and BUT QDB datasets; (ii) We empirically identify Deep SVDD as the superior methodology for this application, demonstrating that it offers the optimal balance between computational efficiency and detection performance; (iii) We validate the efficacy of an integrated classifier-filter system in a realistic deployment simulation incorporating unseen CVD classes and calibrated real-world ambulatory noise. We show that the upstream UAD filter significantly enhances diagnostic robustness, yielding an accuracy improvement of up to 21.0 percentage points over a classifier-only baseline.
@@ II Related Work
The deployment of ML models for ECG analysis in real-world clinical settings introduces significant challenges beyond controlled environments. In this section we review existing methods addressing the challenges most relevant to our work.
The reliability of automated ECG analysis is heavily dependent on input signal quality, as noise artifacts (e.g., muscle activity, electrode motion) prevalent in ambulatory monitoring can severely impair diagnostic accuracy. The literature addresses this through Signal Quality Assessment (SQA) and Uncertainty Quantification (UQ). SQA methods explicitly classify noise types [3] while UQ indirectly flags low-confidence noisy inputs, for example through predictive entropy [4]. Unlike these, our UAD explicitly rejects unsuitable signals regardless of noise type.
@@
UAD methods identify deviations from a learned distribution of normal data without labeled anomalies. In ECG analysis, UAD has been explored primarily to differentiate abnormal patterns from normal sinus rhythm or to enhance supervised learning. Atamny et al. [5] conducted a comparative study of various unsupervised models, including AEs, VAEs, diffusion models, NFs, and Gaussian Mixture Models (GMMs) to differentiate abnormal ECG signals from healthy ones. In a different approach, Jiang et al. [6] employed self-supervised anomaly detection as a pretraining mechanism to improve generalization to rare classes. In contrast, our work concentrates on deploying UAD as an independent, upstream filter designed to reject entirely unseen disease classes and signals rendered unsuitable for further analysis by noise.
@@ III-A Datasets and Preprocessing
We used three datasets in this study. PTB-XL is a large, publicly accessible clinical ECG database comprising 21,799 12-lead ECG recordings, each 10 seconds in duration, from 18,869 patients [10]. Diagnostic labels include five superclasses: Normal (NORM), Myocardial Infarction (MI), Conduction Disturbance (CD), ST/T Change (STTC), and Hypertrophy (HYP). There are potentially multiple superclasses per record, as they are not mutually exclusive. We use the recommended stratified 10-fold cross-validation splits and the 500 Hz ECG signals. The Brno University of Technology ECG Quality Database (BUT QDB) is designed for evaluating signal quality algorithms [11]. It contains 18 long-term, single-lead ECG recordings (sampled at 1,000 Hz) collected during normal daily activities. Expert annotations classify signal quality into three categories: Class 1 (clear visibility of QRS complex, P waves and T waves), Class 2 (only QRS complexes reliably detectable), and Class 3 (signal unsuitable for analysis, QRS complexes undetectable). The MIT-BIH Noise Stress Test database provides realistic noise recordings characteristic of ambulatory ECGs [12]. It includes three half-hour recordings, among which is the electrode motion artifact (’em’) record, utilized in this study for noise injection, that contains significant electrode motion artifacts along with baseline wander and muscle noise. While PTB-XL is a clinical resting dataset, it contains a diverse set of pathologies. By injecting calibrated noise from ambulatory ECGs into the clinical signals, we synthesize the signal degradation characteristic of wearable environments while retaining the ground-truth pathological labels required for rigorous OOD evaluation. All ECG recordings were segmented into 10-second windows. These windows were resampled to a uniform length of 512 timesteps, a configuration determined during pre-screening tests to optimize the balance between computational efficiency and model performance. For the PTB-XL dataset, z-score normalization was applied, as it yielded superior performance. For the BUT QDB dataset, instance normalization proved more effective.
@@ III-B Unsupervised Anomaly Detection Methods
We investigated six UAD methods for detecting anomalous ECG signals. To optimize hyperparameters and evaluate the performance-efficiency trade-off, a NAS using random search (100 trials) was conducted for each method. Models were constrained to a maximum of 512k parameters to ensure suitability for resource-constrained microcontrollers. While a floating-point 32-bit model of this size ($\sim$2MB) would saturate the flash memory of typical low-power MCUs (e.g., STM32U3 [13], NHS52S04 [14]), this search space allows us to accomodate models deployed with 8-bit quantization (reducing footprint to $\sim$0.5MB) and to identify the point of diminishing returns where added complexity no longer yields detection gains. Deep SVDD aims to map normal data into a compact hypersphere of minimum volume [15]. The anomaly score is the squared Euclidean distance to the center in the output space. A 1D ResNet architecture was employed. The NAS optimized the number of layers (2-8), filters (4-1024), kernel sizes (3-11), strides (1-2), and the latent dimension size (8-1024).
@@
MAD utilizes a self-supervised approach similar to masked autoencoders [16]. During training, 5% of time steps (determined via pre-screening) are masked with random values, and the model reconstructs the original values using bidirectional context. The anomaly score is the sum of reconstruction errors calculated by sequentially masking and reconstructing each time step during inference. A standard transformer architecture with an initial patching layer was used. The NAS varied the number of transformer blocks (1-10), hidden dimensions (32-512), attention heads (1-32), linear layer sizes (32-2048), patch size (4-64), and patch overlap (0-75%).
@@
For traditional reconstruction-based methods we used AEs and VAEs to reconstruct given inputs. We utilized the L2 reconstruction error as the anomaly score. Stacked 1D convolution layers were used for both encoder and decoder architectures. The NAS optimized the number of layers (2-8), filters (4-1024), kernel sizes (3-11), strides (1-2), and latent dimension size (12-1024). For the VAE, the KL loss weight was also varied (0.1-1.0).
@@
To use DDPMs for anomaly detection, we adopted the approach of Wyatt et al. [17]: an anomalous input is partially diffused and then reconstructed. The model, trained only on the normal data manifold, attempts to ”repair” abnormal regions during denoising. The reconstruction error serves as the anomaly score. A 1D U-Net architecture with convolution and attention layers was implemented using the denoising-diffusion-pytorch package [18]. The NAS varied the diffusion steps for the noising process (100-1000) and denoising process (10-1000), U-Net architecture parameters (layers (1-10), initial filter (12-256), filter expansion factor (1-8), attention heads (2-8), hidden dimensions (12-256)), and the U-Net objective (predicting noise, the original input or a variable from angular parametrization as in Salimans and Ho [19]).
@@
NFs are exact-likelihood generative models that learn complex data distributions by transforming a simple base distribution through invertible mappings. We implemented a multiscale GLOW architecture [20] with affine coupling layers and invertible 1x1 convolutions, utilizing the normflows package [21]. The exact probability density of a sample serves as the anomaly score. Stacked 1D convolutions were used for the subnetwork within the coupling layers. The NAS optimized the subnetwork architecture (layers (1-5), filters (4-1024), kernel sizes (3-11)) and the flow architecture (number of levels (1-5), coupling blocks (1-10) , split ratios (0.25-0.75) and squeeze factors (1-4)).